### v4

# Distributed Continuous-Time Control via System Level Synthesis

Abstract
This paper designs $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ distributed controllers with local communication and local disturbance rejection. We propose a two-step procedure: first, select closed-loop poles; then, optimize over parameterized controllers. We build on the system level synthesis (SLS) parameterization — primarily used in the discrete-time setting — and extend it to the general continuous-time setting. We verify our approach in simulation on a 9-node grid governed by linearized swing equations, where our distributed controllers achieve performance comparable to that of optimal centralized controllers while facilitating local disturbance rejection.
I Introduction
When regulating the dynamics of large-scale systems (e.g., power grids, process plants), the use of centralized controllers is inefficient due to communication and computational requirements. Distributed controllers offer a more practical alternative; specifically, in scenarios where networked systems which are controlled by networks of sub-controllers, local communication and local disturbance rejection are desirable properties [1]. By local communication, we mean that sub-controllers communicate only with nearby sub-controllers; by local disturbance rejection, we mean that disturbance effects are confined to local neighborhoods. Rigorous definitions are provided in Section II.
Controllers with local communication can be designed using techniques from structured control [12]; however, additional assumptions (e.g., decentralized plants) are often required. Local disturbance rejection is typically induced through tuning or analyzed post-hoc [11]; it is rarely enforced (i.e., via constraints). To the best of our knowledge, no prior work provides a design method for general continuous-time plants that simultaneously accommodates local communication and explicitly enforces disturbance rejection — the goal of this paper is to propose such a method. We leverage the system level synthesis (SLS) parameterization [1] to do so.
The SLS method parameterizes structured controllers using their closed-loop responses. This allows the incorporation of both local communication and local disturbance rejection via sparsity constraints; the resulting problem is convex. Nearly all results on SLS are in the discrete-time setting. The only exception is [6], which is restricted to spatially-invariant and spatially-varying plants with decoupled dynamics.
We first provide the problem setup (Section II) and port the SLS parameterization to continuous time (Section III). We then propose a two-step design procedure: first, select closed-loop poles; then, optimize over closed-loop responses. For the first step, we use simple pole approximation (Section IV); for the second step, we formulate state feedback $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ control as convex optimizations (Section V), providing both LMI- and non-LMI based formulations. Finally, we validate our approach in simulations (Section VI).
Notation: Bold upper-case letters (e.g., $\mathbf{\Phi}(s)$) denote matrices of transfer functions; bold lower case letters (e.g., $\mathbf{u}(s)$) denote signals in $s$ domain. $\mathcal{RH}_{\infty}$ is the Hardy space of real, rational proper transfer functions. $\mathbf{\Phi}(s)\in\frac{1}{s}\mathcal{RH}_{\infty}$ iff $s\mathbf{\Phi}(s)\in\mathcal{RH}_{\infty}$. For matrix $A$, $^{\top}, AA^{*}$ and $\overline{A}$ represent its transpose, complex-conjugate transpose, and complex conjugate, respectively. $\mathbf{0}_{n}\in\mathbb{R}^{n}$ is a zero vector; $I_{n}\in\mathbb{R}^{n\times n}$ is an identity matrix; $\mathbb{D}$ represents the open unit disk.
II Problem setup
II-A Local communication and control
Consider the linear time invariant (LTI) system:
$\dot{x}(t)=Ax(t)+Bu(t)+w(t)$ (1)
where $x(t)\in\mathbb{R}^{n}$, $u(t)\in\mathbb{R}^{m}$, and $w(t)\in\mathbb{R}^{n}$ represent state, control action, and exogenous disturbance. The system contains $N$ interconnected subsystems, each having one or more states. State, control, and disturbance are partitioned into $x]_{i}$, $u]_{i}$, and $w]_{i}$ for each subsystem $i$; system matrices $A$ and $B$ are partitioned into $A]_{ij}$, $B]_{ij}$, which capture subsystem-level interaction. System topology is described by an unweighted directed graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, where vertex $v_{i}$ corresponds to subsystem $i$, and edge $e_{ij}\in\mathcal{E}$ exists whenever $A]_{ij}\neq 0$ or $B]_{ij}\neq 0$. Let $\mathcal{A}\in\{0,1\}^{N\times N}$ be the adjacency matrix for this graph, where by convention $\mathcal{A}_{ii}=1$.
We seek a state feedback controller of the form
$u(t)=\mathcal{K}(x(0:t))$ (2)
where $\mathcal{K}$ is linear and causal. Equivalently, in signal domain:
$\mathbf{u}(s)=\mathbf{K}(s)\mathbf{x}(s)$ (3)
where $\mathbf{K}(s)$ can be implemented as a distributed controller by partitioning along subsystems; row $\mathbf{K}(s)]_{i,;}$ is the sub-controller at subsystem $i$, and $\mathbf{K}(s)]_{ij}$ represents information required by sub-controller $i$ from sub-controller $j$.
Structured communication. We say that controller $\mathbf{K}$ obeys communication structure $\mathcal{S}^{\text{comm}}\in\{0,1\}^{N\times N}$ if $\forall i,j$ s.t. $\mathcal{S}^{\text{comm}}_{ij}=0$, $\mathbf{K}(s)]_{ij}\equiv 0$. To enforce local communication, it suffices to set $\mathcal{S}^{\text{comm}}=\mathcal{A}^{d}$ for some locality parameter $d\in\mathbb{Z}_{+}$, e.g., if $d=1$, then sub-controllers can only communicate if their subsystems are 1-hop neighbors on graph $\mathcal{G}$. An example is given in Fig. 1 (Right).
Structured disturbance rejection. We say that controller $\mathbf{K}$ achieves disturbance rejection with structure $\mathcal{S}^{\text{dist}}\in\{0,1\}^{N\times N}$ if, in the closed-loop system (i.e., system (1) with controller (2)), disturbances $w]_{j}$ can propagate to state $x]_{i}$ only if $\mathcal{S}^{\text{dist}}\neq 0$. To enforce local disturbance rejection, it suffices to set $\mathcal{S}^{\text{dist}}=\mathcal{A}^{d}$, e.g., if $d=1$, then disturbances entering into one subsystem may only spread to another subsystem if they are 1-hop neighbors on graph $\mathcal{G}$.
We now present a clarifying example. Consider a plant with 3 subsystems, with state partition $x]_{1}=\begin{bmatrix}x_{1}&x_{2}\end{bmatrix}^{\top}$, $x]_{2}=\begin{bmatrix}x_{3}&x_{4}\end{bmatrix}^{\top}$, $x]_{3}=x_{3}$ and input partition $u]_{i}=u_{i}$. Then, system matrices $A$ and $B$ can be partitioned as shown in (4); the corresponding graph $\mathcal{G}$ is shown in Fig. 1 (Left).
$A=\left[\begin{array}[]{rr|rr|r}3&2&0&1&0\\ 1&0&0&2&0\\ \hline\cr 0&0&1&1&1\\ 0&0&2&0&0\\ \hline\cr 0&0&5&2&1\end{array}\right],\quad B=\left[\begin{array}[]{r|r|r}0&0&0\\ 1&0&0\\ \hline\cr 0&0&0\\ 0&1&0\\ \hline\cr 0&0&1\end{array}\right]$ (4)
The adjacency matrix for this system is: $\mathcal{A}=\begin{bmatrix}1&1&0\\ 0&1&1\\ 0&1&1\end{bmatrix}$.
If we set $\mathcal{S}^{\text{comm}}=\mathcal{A}$, then sub-controller 1 cannot receive communication from sub-controller 3, and sub-controllers 2 and 3 cannot receive communication from sub-controller 1. If we set $\mathcal{S}^{\text{dist}}=\mathcal{A}$, then disturbances in subsystem 1 may not propagate to any other subsystem; disturbances in subsystem 3 may not propagate to subsystem 1.
II-B Problem statement
We seek a controller (2) that minimizes some objective $f$ and satisfies the aforementioned structured constraints:
$\min_{\mathcal{K}}$ $f(x(t),u(t))$ (5)
s.t. $\dot{x}(t)=Ax(t)+Bu(t)+w(t),\quad u(t)=\mathcal{K}(x(0:t))$
$\mathcal{K}\text{ internally stabilizing, linear, causal}$
$\mathcal{K}\text{ obeys communication structure }\mathcal{S}^{\text{comm}}$
$\mathcal{K}\text{ achieves dist. rejection with structure }\mathcal{S}^{\text{dist}}$
This problem is generally nonconvex due to requirements on both structured communication and disturbance rejection. We use system level synthesis (SLS) to convexify it.
Figure 1: (Left) Graph topology for (4). (Right) Grid topology for Section VI. The central node communicates with all nodes encircled by the solid red line (if $d=1$) or dashed blue line (if $d=2$).
III System level synthesis
We review core elements of the SLS framework and begin translating them into the continuous-time setting. First, take the Laplace transform of (1) and rearrange to get:
$(sI-A)\mathbf{x}(s)=B\mathbf{u}(s)+\mathbf{w}(s)$ (6)
Plugging in equation (3) gives:
$\mathbf{x}(s)=(sI-A-B_{2}\mathbf{K}(s))^{-1}\mathbf{w}(s)=:\mathbf{\Phi}_{x}(s)\mathbf{w}(s)$ (7a)
$\mathbf{u}(s)=\mathbf{K}(s)\mathbf{\Phi}_{x}(s)\mathbf{w}(s)=:\mathbf{\Phi}_{u}(s)\mathbf{w}(s)$ (7b)
where $\mathbf{\Phi}_{x}(s)$ and $\mathbf{\Phi}_{u}(s)$ are the closed-loop responses from disturbance to state and control action, respectively. Theorem 1 (analogous to Theorem 4.1 in [1]) shows that solving over stabilizing controllers ($\mathbf{K}$) is equivalent to solving over an affine subspace of closed-loop responses ($\mathbf{\Phi}_{x},\mathbf{\Phi}_{u}$).
Figure 2: Controller implementation (9), in closed loop with the plant and exogeneous disturbances/noise. Exogeneous signals $\boldsymbol{\delta}_{u}$ and $\boldsymbol{\delta}_{y}$ are used to verify internal stability of the system.
Theorem 1.
For system (1) evolving under causal, linear state feedback controller (3), the following hold:
Statement 1. The affine subspace [^1] This subspace is nontrivial if system $(A,B)$ is stabilizable. defined by:
$\begin{bmatrix}sI-A&-B\end{bmatrix}\begin{bmatrix}\mathbf{\Phi}_{x}(s)\\ \mathbf{\Phi}_{u}(s)\end{bmatrix}=I$ (8a)
$\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\in\frac{1}{s}\mathcal{RH}_{\infty}$ (8b)
parameterizes all closed-loop responses $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$ that are achievable by an internally stabilizing controller $\mathbf{K}(s)$.
Statement 2. $\forall\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$ satisfying (8), the causal linear controller $\mathbf{K}(s)=\mathbf{\Phi}_{u}(s)\mathbf{\Phi}_{x}(s)^{-1}$, implemented in feedback configuration (9) (see Fig. 2) is internally stabilizing and achieves closed-loop responses $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$.
$\begin{split}\mathbf{v}(s)&=\mathbf{x}(s)+(I-s\mathbf{\Phi}_{x}(s))\mathbf{v}(s)\\ \mathbf{u}(s)&=s\mathbf{\Phi}_{u}(s)\mathbf{v}(s)\end{split}$ (9)
Proof.
(Statement 1) Any stabilizing controller $\mathbf{K}(s)$ has associated closed-loop responses $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$ that satisfy (8); plug (7) into (8) for the desired result.
(Statements 1 & 2) We now show that if some transfer function matrices $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$ satisfy (8), there exists an internally stabilizing controller which achieves closed-loop responses $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$. It suffices to show that the controller described by Statement 2 fulfils these requirements.
First, we prove that $\mathbf{\Phi}_{x}(s)$ is invertible. We know that $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$ satisfy (8); rearranging (8a) yields:
$\mathbf{\Phi}_{x}(s)=(sI-A)^{-1}(I+B\mathbf{\Phi}_{u}(s))$
So $\mathbf{\Phi}_{x}(s)$ being invertible is equivalent to $I+B\mathbf{\Phi}_{u}(s)$ being invertible which, in turn, is equivalent to $I+B\mathbf{\Phi}_{u}(\infty)$ being invertible [14]. Since $\mathbf{\Phi}_{u}(s)$ is strictly proper by (8), $\mathbf{\Phi}_{u}(\infty)=0$ and $I+B\mathbf{\Phi}_{u}(\infty)=I$, which is invertible.
Next, we show that this controller achieves the desired response. Plug $\mathbf{K}(s)=\mathbf{\Phi}_{u}(s)\mathbf{\Phi}_{x}(s)^{-1}$ into (6) and (3) and apply (8a) to yield the desired closed-loop responses $\mathbf{x}(s)=\mathbf{\Phi}_{x}(s)\mathbf{w}(s)$ and $\mathbf{u}(s)=\mathbf{\Phi}_{u}(s)\mathbf{w}(s)$. Finally, to show that this controller realization is internally stabilizing, consider perturbations $(\mathbf{w},\boldsymbol{\delta}_{u},\boldsymbol{\delta}_{y})$ to the plant, control input, and plant output, respectively (see Fig. 2). Build nine transfer functions from $(\mathbf{w},\boldsymbol{\delta}_{u},\boldsymbol{\delta}_{y})$ to $(\mathbf{x},\mathbf{u},\mathbf{v})$ and verify that all are stable; follow the procedure from Section 4 of [1]. ∎
In this parameterization, structured (and therefore local) communication and disturbance rejection can be achieved by enforcing sparsity constraints on closed-loop responses:
$\mathbf{\Phi}_{x}(s)\in\mathcal{S}_{x},\quad\mathbf{\Phi}_{u}(s)\in\mathcal{S}_{u}$ (10)
where $\mathcal{S}$ represents sparsity constraints encoding structured communication and disturbance rejection. With respect to the original problem (5), $\mathcal{S}=\mathcal{S}^{\text{comm}}\cap\mathcal{S}^{\text{dist}}$ [^2] One limitation is that in this framework, it is impossible to enforce structured communication without enforcing structured disturbance rejection, and vice versa; see [8] for a strategy that mitigates this.
Matrices $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$ are also used to implement the controller via (9), which allows us to avoid inversion of $\mathbf{\Phi}_{x}$ and thus, preserve sparsity. We can then rewrite the original problem (5) in the SLS parameterization as:
[…]
where $g(\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s))$ is formulated to be equivalent to the original control objective (see Section V for examples). If objective $g$ is convex, then this problem can be solved using convex optimization methods, since both constraints are affine with respect to the decision variables.
Remark: Enforcing sparsity constraints on the decision variables can also facilitate scalability if the objective and constraints are separable. For example, if minimizing $g$ for each column of $\mathbf{\Phi}_{x}$ and $\mathbf{\Phi}_{u}$ is equivalent to minimizing $g$ over the full matrices, then instead of solving optimization (11) directly, we can split it into smaller column-wise optimization problems be solved in parallel. If the decision variables are constrained to be sparse, the size of each column-wise problem will remain independent of size of the overall network. See [1, 7] for additional discussion.
IV Closed-loop pole selection
IV-A Finite-dimensional truncation
Problem (11) involves solving over the infinite-dimensional affine subspace (8). In previous work on discrete-time SLS, finite impulse response approximations are used to truncate this problem into a finite-dimensional one; we restrict ourselves to closed-loop responses that satisfy (8) and place all closed-loop poles at zero, up to some fixed multiplicity. For continuous-time systems, an analogous approach is to first select some fixed, finite multiset $\mathcal{P}$ of closed-loop poles, and restrict ourselves to solving over all closed-loop responses that satisfy (8) and place closed-loop poles at $\mathcal{P}$. This is related to the notion of Ritz approximations in continuous time, which utilize a single pole with varying multiplicities [10], as well as simple pole approximation in discrete-time [4].
Given $\mathcal{P}$, some multiset of poles which are contained in the open left-half-plane (OLHP), we restrict ourselves to considering closed-loop responses of the form:
$\mathbf{\Phi}_{x}(s)$ $=\sum_{l=1}^{K}\sum_{j=1}^{m_{l}}\Phi_{x}(l,j)\frac{1}{{(s-p_{l})}^{j}}$ (12)
$\mathbf{\Phi}_{u}(s)$ $=\sum_{l=1}^{K}\sum_{j=1}^{m_{l}}\Phi_{u}(l,j)\frac{1}{{(s-p_{l})}^{j}}$
where $K$ is the number of distinct poles in $\mathcal{P}$, and $m_{l}$ is the number of times $p_{l}$ appears in $\mathcal{P}$. We will refer to $\Phi_{x}(l,j)$ and $\Phi_{u}(l,j)$ as the spectral components of $\mathbf{\Phi}_{x}(s)$ and $\mathbf{\Phi}_{u}(s)$, respectively. Now, we can rewrite constraint (8) in terms of the spectral components of the closed-loop responses:
Lemma 1.
Consider closed-loop responses of the form (12). If they are real-valued, then they lie in the affine subspace defined by (8) if and only if the following hold:
[…]
Proof.
(8b) holds by construction of $\mathcal{P}$ and (12). To show that (8a) holds, plug (12) into (8a) and write out the constraint at each index (similar to [3]). ∎
Remark: The number of poles required to satisfy (13) does not depend on system size. For example, consider system (1) with an arbitrary number of states and full rank $B$ matrix. It is sufficient to choose a single pole $p<0$; then, $\mathbf{\Phi}_{x}(s)=\frac{1}{s-p}I_{n}$ and $\mathbf{\Phi}_{u}(s)=\frac{1}{s-p}B^{-1}$ satisfy (13).
Remark: In general, spectral components $\Phi_{x}(l,j)$, $\Phi_{u}(l,j)$ need not be real; see Section V for details.
We use simple pole approximation (SPA), where all poles have multiplicity 1 (i.e., $m_{l}=1$ $\forall$ $l$). Then, (12) becomes:
$\mathbf{\Phi}_{x}(s)=\sum_{l=1}^{K}\Phi_{x}(l)\frac{1}{(s-p_{l})},\hskip 5.0pt\mathbf{\Phi}_{u}(s)=\sum_{l=1}^{K}\Phi_{u}(l)\frac{1}{(s-p_{l})}$ (14)
and constraint (13) becomes:
$\sum_{l=1}^{K}\Phi_{x}(l)=I,\quad B\Phi_{u}(l)=(p_{l}I-A)\Phi_{x}(l)$ (15)
The remainder of this section focuses on SPA. However, optimization programs provided in Section V can be used with poles from SPA or other methods (e.g., Ritz). We briefly compare SPA against other pole selection methods in Section VI and defer detailed exploration to future work.
IV-B Simple pole approximation
We now present SPA for continuous-time SLS. Our approach is inspired by [4]; here, we devise a continuous-time version and rigorously demonstrate (Theorems 3-4, Lemma 2) that poles generated via SPA indeed let us approximate the optimal continuous-time closed-loop responses.
Let $\mathcal{P}$ now denote a set of $K$ poles generated by SPA, where $K$ is even. We first generate discrete-time poles $z_{k}\in\mathbb{D}$ using the Archimedes spiral and its reflection, then map them into the OLHP using the bilinear transform:
$\text{for }k=1\ldots\frac{K}{2},\quad\theta_{k}=2\sqrt{\pi k},\quad r_{k}=\sqrt{\frac{k}{K+1}}$ (16)
$z_{k}=r_{k}e^{j\theta_{k}},\quad z_{-k}=r_{k}e^{-j\theta_{k}},\quad p_{k}=\frac{z_{k}-1}{z_{k}+1}$
To show that this approximation method approaches the optimal solution, we first recall a result from prior work.
Theorem 2.
(Theorem 4 in [4]) Let $\mathcal{Z}_{\mathcal{P}}$ be a set of $K$ poles generated by discrete-time SPA, i.e., $\mathcal{Z}_{\mathcal{P}}:=\{z_{k}\}_{k=-0.5K}^{0.5K}$. Let $\mathcal{Z}_{\mathcal{Q}}$ be some arbitrary set of poles in $\mathbb{D}$. Define $D(\mathcal{Z}_{\mathcal{P}})$ as the max-min distance between poles in $\mathcal{Z}_{\mathcal{P}}$ and $\mathcal{Z}_{\mathcal{Q}}$, i.e.,
$D(\mathcal{Z}_{\mathcal{P}}):=\underset{x\in\mathcal{Z}_{\mathcal{Q}}}{\max}(\underset{y\in\mathcal{Z}_{\mathcal{P}}}{\min}|x-y|)\\$ (17)
Then, $\underset{K\to+\infty}{\lim}D(\mathcal{Z}_{\mathcal{P}})\to 0$.
This theorem shows that SPA-generated poles can approximate the poles ($\mathcal{Z}_{\mathcal{Q}}$) of any discrete-time transfer function matrix. We begin to extend this to continuous-time: first, we show that poles close in the unit disk will also be close in the OLHP after applying the bilinear transform.
Theorem 3.
Let $\mathcal{P}$ and $\mathcal{Q}$ be two sets of poles in the OLHP. Consider $p\in\mathcal{P}$, $q\in\mathcal{Q}$, $z_{p}=\frac{1+p}{1-p}$ and $z_{q}=\frac{1+q}{1-q}$. Define the following constants:
$c_{1}$ $=\frac{2c_{4}}{c_{3}(c_{3}+c_{4})},$ $\quad c_{2}$ $=\frac{c_{3}(c_{3}+c_{4})}{2}$
$c_{3}$ $=\max_{q\in\mathcal{Q}}|1-q|,$ $\quad c_{4}$ $=1+\min_{q\in\mathcal{Q}}|q|$
If $|z_{p}-z_{q}|\leq c_{1}$, then the following must hold:
$|p-q|\leq c_{2}|z_{p}-z_{q}|,\quad|p-q|\leq c_{4}$
Proof.
Since $p=\frac{z_{p}-1}{z_{p}+1}$ and $q=\frac{z_{q}-1}{z_{q}+1}$, we have:
$|p-q|=\frac{2}{|1+z_{p}||1+z_{q}|}|z_{p}-z_{q}|$ (18)
One of the right-hand-side terms can be upper-bounded:
$\begin{split}\frac{2}{|1+z_{p}||1+z_{q}|}&=\frac{2}{|(1+z_{q})+(z_{p}-z_{q})||1+z_{q}|}\\ &\leq\frac{2}{||1+z_{q}|-|z_{p}-z_{q}|||1+z_{q}|}\end{split}$ (19)
We can further refine this bound as follows: since $|1+z_{q}|=\frac{2}{|1-q|}$, we have $|1+z_{q}|\geq\frac{2}{{\max}_{q\in\mathcal{Q}}|1-q|}=\frac{2}{c_{3}}$. Since $|z_{p}-z_{q}|\leq c_{1}$, we have that $|z_{p}-z_{q}|\leq\frac{c_{4}}{c_{4}+c_{3}}\frac{2}{c_{3}}\leq\frac{c_{4}}{c_{4}+c_{3}}|1+z_{q}|<|1+z_{q}|$. Then, we can rewrite (19) as:
$\frac{2}{|1+z_{p}||1+z_{q}|}$ $\leq\frac{2}{(|1+z_{q}|-|z_{p}-z_{q}|)|1+z_{q}|}$ (20)
$\leq\frac{2}{(|1+z_{q}|-\frac{c_{4}}{c_{4}+c_{3}}|1+z_{q}|)|1+z_{q}|}$
$=\frac{2(c_{3}+c_{4})}{c_{3}|1+z_{q}|^{2}}\leq c_{2}$
Plugging this into (18) yields $|p-q|\leq c_{2}|z_{p}-z_{q}|$. Combining this with $|z_{p}-z_{q}|\leq c_{1}$ yields $|p-q|\leq c_{1}c_{2}=c_{4}$. ∎
This leads to a continuous-time analogue of Theorem 2:
Lemma 2.
Let $\mathcal{P}$ be a set of $K$ poles generated by continuous-time SPA as described in equation (16). Let $\mathcal{Q}$ be some arbitrary set of poles in the OLHP. Define $D(\mathcal{P})$ as the max-min distance between poles in $\mathcal{P}$ and $\mathcal{Q}$, i.e.,
$D(\mathcal{P})=\underset{x\in\mathcal{Q}}{\max}(\underset{y\in\mathcal{P}}{\min}|x-y|)$ (21)
It is always possible to select $K$ such that $D(\mathcal{P})\leq c_{4}$.
Proof.
Let $\mathcal{Z}_{\mathcal{P}}$ and $\mathcal{Z}_{\mathcal{Q}}$ represent the bilinear transforms of sets $\mathcal{P}$ and $\mathcal{Q}$. For any $z_{p}\in\mathcal{Z}_{\mathcal{P}}$ and $z_{q}\in\mathcal{Z}_{\mathcal{Q}}$, $|z_{p}-z_{q}|\leq D(\mathcal{Z}_{\mathcal{P}})$ by definition. From Theorem (2), we can always choose $K$ to make $D(\mathcal{Z}_{\mathcal{P}})$ arbitrarily small; thus, we can always choose $K$ so that $|z_{p}-z_{q}|\leq c_{1}$. Applying Theorem (3) gives the desired result. ∎
Now, the final result: transfer function matrices using SPA-generated poles can approximate arbitrary stable transfer function matrices in continuous time. We write transfer function matrices with SPA-generated poles as:
$\mathbf{G}(s)=\sum_{p\in\mathcal{P}}G_{p}\frac{1}{s-p}$ (22)
where spectral components $\{G_{p}\}_{p\in\mathcal{P}}$ are constant matrices.
Theorem 4.
Let $\mathcal{P}$ be a set of poles generated by continuous-time SPA per (16); select $K$ so that $D(\mathcal{P})\leq c_{4}$. Then, for any transfer function matrix $\mathbf{S}(s)\in\frac{1}{s}\mathcal{RH}_{\infty}$, there exists constants $c_{s},c_{s}^{\prime}>0$ and matrices $\{G_{p}\}_{p\in\mathcal{P}}$ such that
$\left\|\sum_{p\in\mathcal{P}}G_{p}\frac{1}{s-p}-\mathbf{S}(s)\right\|_{\mathcal{H}_{2}}\leq{c_{s}}D(\mathcal{P})$ (23a)
$\left\|\sum_{p\in\mathcal{P}}G_{p}\frac{1}{s-p}-\mathbf{S}(s)\right\|_{\mathcal{H}_{\infty}}\leq{c_{s}^{\prime}}D(\mathcal{P})$ (23b)
This result and its proof are analogous to Theorem 1 in [4]; a proof sketch is provided in the Appendix.
V $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ Control via SLS
SLS problem (11) is infinite-dimensional. To bypass this, we constrain closed-loop responses $\mathbf{\Phi}_{x}$ and $\mathbf{\Phi}_{u}$ to have pre-generated stable poles, and optimize over spectral components. These results can be used with poles generated through SPA or other methods (e.g., Ritz).
To simplify notation, we re-order the SPA-generated pole set $\mathcal{P}$: $\mathcal{P}=\{p_{1},p_{-1},\dots p_{\frac{K}{2}},p_{-\frac{K}{2}}\}$. For the remainder of the paper (except the Appendix), we re-label these poles as $\mathcal{P}=\{p_{1},p_{2},\dots p_{K}\}$, with the convention that $(p_{2i-1},p_{2i})$ is a complex conjugate pair. While poles can be complex, transfer function matrices $\mathbf{\Phi}_{x}(s)$ and $\mathbf{\Phi}_{u}(s)$ must be real-valued; thus, spectral components must satisfy the following:
$\quad\forall i\in\{1,2,\ldots,\frac{K}{2}\},\quad\Phi_{x}(2i-1)$ $=\overline{\Phi_{x}(2i)},$ (24)
$\Phi_{u}(2i-1)$ $=\overline{\Phi_{u}(2i)},$
Then, for poles chosen via SPA, feasibility constraints (8) can be enforced via (24) and (15). Similarly, sparsity constraints (10) can be enforced via:
$\Phi_{x}(l)\in\mathcal{S}_{x},\quad\Phi_{u}(l)\in\mathcal{S}_{u}$ (25)
The SLS-based control problem (11) now becomes:
$\min_{\Phi_{x}(l),\ \Phi_{u}(l),\text{ }l=1\ldots K}$ ${g(\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s))}$ (26)
[…]
where we optimize over spectral components. We now derive expressions for objective $g$ in terms of spectral components.
V-A State-space realization
We consider linear quadratic (LQ) cost formulations with state penalty matrix $Q\succeq 0$ and input penalty matrix $R\succ 0$. We introduce augmented variable $\mathbf{\Psi}(s)$:
$\mathbf{\Psi}(s):=\begin{bmatrix}Q^{\frac{1}{2}}&0\\ 0&R^{\frac{1}{2}}\end{bmatrix}\begin{bmatrix}\mathbf{\Phi}_{x}(s)\\ \mathbf{\Phi}_{u}(s)\end{bmatrix}$ (27)
We can write the spectral components of $\mathbf{\Psi}(s)$ as linear combinations of spectral components of $\mathbf{\Phi}_{x}(s)$ and $\mathbf{\Phi}_{u}(s)$:
$\mathbf{\Psi}(s)$ $=\sum_{l=1}^{K}\Psi(l)\frac{1}{s-p_{l}}$ (28)
$\Psi(l)$ $=\begin{bmatrix}Q^{\frac{1}{2}}&0\\ 0&R^{\frac{1}{2}}\end{bmatrix}\begin{bmatrix}\Phi_{x}(l)\\ \Phi_{u}(l)\end{bmatrix}$
The objective is to minimize $\|\mathbf{\Psi}(s)\|_{\mathcal{H}_{2}}$ for $\mathcal{H}_{2}$ control, and $\|\mathbf{\Psi}(s)\|_{\mathcal{H}_{\infty}}$ for $\mathcal{H}_{\infty}$ control. To facilitate this, we use the following state-space realization of (28):
$\dot{x}$ $=\mathcal{A}x+\mathcal{B}u=:\begin{bmatrix}p_{1}I_{n}&&\\ &\ddots&\\ &&p_{K}I_{n}\end{bmatrix}x+\begin{bmatrix}I_{n}\\ \vdots\\ I_{n}\end{bmatrix}u$
$y$ $=\mathcal{C}x=:\begin{bmatrix}\Psi(1)&\ldots&\Psi(K)\end{bmatrix}x$ (29)
When using the SPA-generated poles, all poles are complex, which means realization (V-A) is complex. However, if (24) is satisfied, then we can construct orthogonal matrix
$T=\text{blkdiag}(U,U,\ldots U),\quad U=\begin{bmatrix}\frac{1+j}{2}&\frac{1-j}{2}\\ \frac{1-j}{2}&\frac{1+j}{2}\end{bmatrix}$ (30)
such that the similarity-transformed matrices
$\tilde{\mathcal{A}}=T^{*}\mathcal{A}T,\quad\tilde{\mathcal{B}}=T^{*}\mathcal{B},\quad\tilde{\mathcal{C}}=\mathcal{C}T$ (31)
are real-valued. Even when poles are not generated by SPA, as long as $\mathbf{\Psi}(s)$ is real-valued, we can always find a real-valued state-space realization for the controller. For general non-repeated poles, partition $\mathcal{A}$ into $\mathcal{A}=\text{blkdiag}(\mathcal{A}_{c},\mathcal{A}_{r})$, where $\mathcal{A}_{c}$ contains complex conjugate poles (where adjacent values are conjugate pairs, similar to SPA) and $\mathcal{A}_{r}$ contains real-valued poles. Then, apply $T=\text{blkdiag}(U,U,\ldots U,I)$. For repeated poles (e.g., Ritz), let the $(i,j)^{\text{th}}$ element of transfer function matrix $\mathbf{\Psi}(s)$ be written as
$\mathbf{\Psi}_{ij}(s)=\sum_{l=1}^{K}\frac{\Psi_{ij}(l)}{(s-p)^{l}}$ (32)
A state-space realization for $\mathbf{\Psi}_{ij}(s)$ is:
$\mathcal{A}^{ij}$ $=\begin{bmatrix}p&1&0&\cdots&0\\ 0&p&1&\cdots&0\\ 0&0&p&\cdots&0\\ \vdots&\vdots&\vdots&&\vdots\\ 0&0&0&\cdots&p\end{bmatrix},\quad\mathcal{B}^{ij}=\begin{bmatrix}1\\ 1\\ \vdots\\ 1\end{bmatrix}$, (33)
$\mathcal{C}^{ij}$ $=\begin{bmatrix}\Psi_{ij}(K)&\mathcal{C}^{ij}_{K-1}\ldots\mathcal{C}^{ij}_{1}\end{bmatrix},\text{where }$
$\mathcal{C}^{ij}_{k}=\Psi_{ij}(k)-\Psi_{ij}(k+1)\text{ for }k=1\ldots K-1$
where $\mathcal{A}^{ij}\in\mathbb{R}^{K\times K}$. Concatenate state-space realizations $(\mathcal{A}^{ij},\mathcal{B}^{ij},\mathcal{C}^{ij})$ to produce a realization for $\mathbf{\Psi}(s)$.
The remainder of this section computes $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ objectives using these state-space realizations.We use ($\tilde{\mathcal{A}},\tilde{\mathcal{B}},\tilde{\mathcal{C}}$) to indicate the appropriate real-valued state-space realization.
V-B $\mathcal{H}_{2}$ control
From linear matrix inequality (LMI) results in [13], we know that $\left\|\mathbf{\Psi}(s)\right\|_{\mathcal{H}_{2}}<\eta$ if $\exists P\succ 0$, $W\succ 0$ such that
$\begin{bmatrix}\tilde{\mathcal{A}}^{\top}P+P\tilde{\mathcal{A}}&P\tilde{\mathcal{B}}\\ \tilde{\mathcal{B}}^{\top}P&-I\end{bmatrix}\prec 0,\quad\begin{bmatrix}P&\tilde{\mathcal{C}}^{\top}\\ \tilde{\mathcal{C}}&W\end{bmatrix}\succ 0,\quad\text{trace}(W)<\eta$ (34)
Then, we can write the $\mathcal{H}_{2}$ problem as:
$\min_{\begin{subarray}{c}\Phi_{x}(l),\ \Phi_{u}(l),\text{ }l=1\ldots K\\ P\succ 0,\ W\succ 0,\ \eta\geq 0\end{subarray}}{\quad\eta}$ (35)
[…]
which is convex. Constraint (34) can be partitioned into smaller parallel constraints by using partitioned matrices (similar idea as (33)); then, this problem is separable and therefore scalable (see Section III). When poles are distinct (e.g., generated via SPA), we can equivalently write:
$\min_{\Phi_{x}(l),\ \Phi_{u}(l),\text{ }l=1\ldots K}$ ${\sum_{i=1}^{m}\sum_{j=1}^{n}R_{ij}}$ (36)
[…]
$R_{ij}=\sum_{l=1}^{K}\lim_{s\to p_{l}}(s-p_{l})\overline{\Psi_{ij}(s)}\Psi_{ij}(s)$
V-C $\mathcal{H}_{\infty}$ control
First, recall the KYP lemma (Lemma 4.1 in [5]):
Lemma 3.
Consider a strictly proper transfer function $G(s)$ with real-valued realization $(\tilde{\mathcal{A}},\tilde{\mathcal{B}},\tilde{\mathcal{C}},0)$. Then, the statement
[…]
is equivalent to the statement
$\exists X\succ 0\text{ s.t.}\begin{bmatrix}{\tilde{\mathcal{A}}^{\top}}X+X\tilde{\mathcal{A}}&X\tilde{\mathcal{B}}&\tilde{\mathcal{C}}^{\top}\\ {\tilde{\mathcal{B}}^{\top}}X&-\gamma I&0\\ \tilde{\mathcal{C}}&0&-\gamma I\end{bmatrix}\prec 0$ (38)
Using this, we can write the convex $\mathcal{H}_{\infty}$ program:
$\min_{\begin{subarray}{c}\Phi_{x}(l),\ \Phi_{u}(l),\text{ }l=1\ldots K\\ X\succ 0,\ \gamma\geq 0\end{subarray}}{\quad\eta}$ (39)
[…]
VI Simulations
We simulate our controller synthesis technique on two plants, setting $Q$ and $R$ as identity matrices. Simulations can be reproduced using the SLS-MATLAB toolbox [9].
VI-A Chain plant
This unstable plant has 11 nodes; $A$ is tri-diagonal, with $A_{ii}=0.6$, $A_{i,i-1}=A_{i-1,i}=0.4$, and $B=I$. We use a distributed $\mathcal{H}_{2}$ controller with 4 SPA-generated poles and locality parameter $d=2$.The $\mathcal{H}_{2}$ cost of our controller is $1.18$ times the cost of the centralized LQR controller. By design, this controller achieves local communication and local disturbance rejection (see Fig. 3).
Figure 3: Chain plant with SLS controller. A disturbance is injected at node 6 at time $t=0$; it propagates only to nodes 4, 5, 6, 7, and 8 (i.e., nodes within $d=2$ distance of node 6) before being rejected.
VI-B Grid plant
We consider linearized swing equations, as is done in [1]. The state of node $i$ is $x]_{i}:={[\theta_{i}\quad\dot{\theta}_{i}]}^{\top}$, with dynamics
[…]
$+\begin{bmatrix}0\\ 1\end{bmatrix}u_{i}+\begin{bmatrix}0\\ 1\end{bmatrix}w_{i}$
where parameters $m_{i}^{-1}$, $\beta_{i}$, and $k_{ij}$ represent inertia, damping, and coupling, respectively; $\theta_{i}$, $w_{i}$, and $u_{i}$ are phase angle deviation, disturbance, and control action, respectively; $\mathcal{N}_{i}$ indicates neighbors of node $i$. We generate a random grid topology (Fig. 1) and random parameters using similar distributions as in [1]; the resulting plant is marginally stable.
We design $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ controllers with varying locality parameters ($d$) and number of poles. Results are shown in Fig. 4. As $d$ or the number of poles decreases, performance becomes less optimal; the $\mathcal{H}_{\infty}$ controller performance is less sensitive in this regard than $\mathcal{H}_{2}$. Note that near-optimal $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ controllers can be designed via SPA using only 6 poles, even though the true optimal closed-loop has 18 poles. We also include results using Ritz approximations for $\mathcal{H}_{2}$, varying the multiplicity of the pole instead of the number of poles; we find that this method is impractical as it is quite sensitive to the choice of pole placement.
We also compared our continuous-time controllers to discrete-time SLS controllers with zero-order hold. We let each controller have the same number of free variables, using horizon $T=7$ for discrete-time and 6 SPA-generated poles for continuous-time. The resulting $\mathcal{H}_{2}$ costs are 5.22 for the continuous-time case, and 5.48, 5.32, and 6.98 for the discrete-time case corresponding to sampling time $\Delta t=0.05$, $0.1$, and $0.2$ seconds. The continuous-time controller outperforms all discrete-time controllers; as expected, there is benefit in designing directly in the continuous-time domain.
Figure 4: Grid plant with various SLS controllers. All costs are normalized by the cost of the optimal centralized controller. For this system, $d=5$ is equivalent to a centralized controller since the diameter of the graph is 5.